# OrbitZoo: Multi-Agent Reinforcement Learning Environment for Orbital Dynamics

### v1

@@ Appendix 0.A Orbital Mechanics
This section provides an overview of key orbital mechanics concepts to support a clearer understanding of the concepts presented throughout the paper. We begin by explaining the most commonly used coordinate systems, then discuss how body propagation is usually managed, and conclude with our approach to realistic thrust representation.
0.A.1 Coordinate Systems
@@
Figure 6: Representation of Cartesian position ($r$) and velocity ($\dot{r}$).
Keplerian elements, as Figure 7 shows, describe the orbital motion using six parameters: the semi-major axis $(a)$, eccentricity $(e)$, inclination $(i)$, argument of perigee $(\omega)$, longitude of the ascending node $(\Omega)$, and anomaly (which can be represented in three ways, as seen in Figure 8). However, the mean anomaly is commonly used due to its linear evolution over time.
Figure 7: Representation of Keplerian Orbital Elements. Here, it is assumed that the orbit is perfectly circular ($e=0$), and consequently, the semi-major axis corresponds to the Euclidean distance to the primary focus ($M$).
Figure 8: Types of anomaly. A body $S$ orbits a central body $C$ in an elliptical orbit with semi-major axis $a$. The true anomaly ($\theta$) is the actual angular position of $S$ measured from the periapsis (closest point to $C$). The mean anomaly ($M$) is the angle assuming uniform motion over the same orbital period (represented by body $S^{\prime}$). Eccentric anomaly ($E$) is an intermediate angle used to relate $M$ and $\theta$, representing the position of $S$ on the circular orbit ($Q$).
These elements directly relate to the orbital geometry and are particularly useful for characterizing two-body motion. However, they can become undefined or poorly behaved in special cases, such as circular or equatorial orbits. These scenarios, referred to as singularities, pose significant challenges for numerical integration and optimization problems. Equinoctial elements are therefore preferred for RL, as they avoid these singularities by using five parameters (plus anomaly): $(a,e_{x},e_{y},h_{x},h_{y})$, where $e_{x}$ and $e_{y}$ represent components of the eccentricity vector, and $h_{x}$ and $h_{y}$ describe the inclination vector.
0.A.2 Thrust Representation
@@
where $\times$ represents the cross product.
@@
$\mathbf{T}_{\text{RSW}}=T(\cos{\theta}\hat{s}+\sin{\theta}(\cos{\phi}\hat{r}+\sin{\phi}\hat{w})).$ (2)
@@ 0.A.3 Propagation
$F=-\frac{GMm}{\|r\|^{2}}\hat{r},$ (5)
[…]
@@
Although this model forms the foundation of orbital mechanics and is commonly used in RL, analytical models like Simplified General Perturbations (SGP) incorporate additional perturbative forces – such as atmospheric drag, solar radiation pressure, and Earth’s oblateness – to provide a fast and relatively accurate method of propagating bodies.
Numerical propagators provide a powerful alternative to SGP by solving the equations of motion through numerical integration, enabling precise modeling of complex orbital dynamics. Unlike SGP, numerical propagators can incorporate a wide range of perturbative forces, including higher-order gravitational harmonics, third-body effects, atmospheric drag, and solar radiation pressure. This flexibility allows them to handle scenarios requiring high precision, such as long-term orbit predictions and mission-critical maneuvers.
@@
$f(t,s)=\frac{d}{dt}s=\frac{d}{dt}\begin{bmatrix}r\\ \dot{r}\\ m\end{bmatrix}=\begin{bmatrix}\dot{r}\\ \ddot{r}\\ \dot{m}\end{bmatrix}=\begin{bmatrix}\dot{r}\\ -\frac{\mu_{E}}{\|r\|^{2}}\hat{r}+\frac{T}{m}+a_{\text{env}}\\ -\frac{\|T\|}{I_{\text{sp}}g}\end{bmatrix}.$ (7)
@@
Appendix 0.B Mathematical Definitions
@@
0.B.1 Proximal Policy Optimization (PPO)
The actor loss in PPO is defined based on the ratio $r_{t}(\theta)$, which measures the probability of taking an action $a_{t}$ under the current policy compared to the old policy:
@@ 0.B.1 Proximal Policy Optimization (PPO)
0.B.2 Generalized Advantage Estimation (GAE)
Generalized Advantage Estimation (GAE) provides a flexible way to compute the advantage by balancing bias and variance through the $\lambda$ parameter:
@@ 0.B.2 Generalized Advantage Estimation (GAE)
Appendix 0.C Computational Performance
@@
OrbitZoo imposes no strict limit on the number of bodies in a system, with scalability constrained only by the available hardware (0.C.1). However, it is essential to assess how increasing the number of bodies and introducing complex dynamics impact simulation speed and parallelization efficiency. These aspects are examined in the following subsections (0.C.2 and 0.C.3).
Although not yet implemented, OrbitZoo experiments could benefit from libraries mentioned in 3.4 to accelerate training.
0.C.1 Hardware Specifications
@@ 0.C.2 Scalability
To address the first question, we evaluate system performance when propagating 1, 5, 10, 100, 1000, and 10000 bodies. For the second question, we analyze simulation speed starting with a simple Newtonian gravity model assuming a perfectly spherical central body. We then introduce a more complex gravity model using spherical harmonics (HolmesFeatherstone) and further incorporate perturbative forces, including third-body effects (Sun and Moon), solar radiation pressure, and drag.
@@
Figure 10: Time per step (in seconds) by number of bodies and active forces. Each point represents an average of 100 propagation steps of 10 seconds.
0.C.3 Parallelization
@@
As shown in Figure 1, OrbitZoo integrates with two key external components: PettingZoo and Orekit. PettingZoo provides a framework for sequential or parallelized step computations, while Orekit handles orbital dynamics within each step. Since orbital MARL missions assume all agents act simultaneously rather than in turns, OrbitZoo propagates systems in a parallel manner.
@@
OrbitZoo leverages PettingZoo’s parallelization by implementing the ParallelEnv class. To assess the impact of this choice, we compared simulation times between parallel and sequential step computations. Specifically, we measured how much longer sequential execution takes relative to parallel execution as the number of bodies increases (1, 5, 10, 100, 1000, and 10000 bodies).
Figure 11 shows that using a sequential propagation architecture instead of a parallel one leads to a linear increase in step computation time with the number of bodies. For instance, with 100 bodies, sequential propagation takes nearly 0.1 seconds longer per step than parallel propagation, while with 10000 bodies, the difference exceeds 1 second per step. Given that a single RL episode can consist of hundreds or thousands of steps, this improvement significantly enhances simulation efficiency.
@@
Figure 11: Additional time per step (in seconds) when using sequential computation instead of parallel, by the number of bodies. Each point represents an average of the difference of 100 propagation steps of 10 seconds, between the sequential and parallel step times.
Orekit also offers tools for parallelizing propagations, such as the PropagatorsParallelizer class, which creates new threads to speed up computations. OrbitZoo could benefit from these tools to further improve simulation time. However, this has not been implemented yet, as parallelization in Java (native to Orekit) is handled differently in the Python wrapper. This discrepancy presents a challenge, but it will be an area of future research. As of now, the individual propagation of bodies remains the primary bottleneck for parallelization and scalability within OrbitZoo.
Appendix 0.D Experiments
To evaluate OrbitZoo’s capabilities, we designed a series of missions utilizing a state-of-the-art algorithm (0.D.1). The first two missions (0.D.2, 0.D.3) focus on orbital transfers (OTs), where agents learn strategies to reach static and dynamic targets, respectively, while accounting for real-world factors such as mass loss due to fuel consumption. The third mission (0.D.4) involves a satellite learning to minimize collision probability with debris in the days leading up to the predicted closest approach, all while maintaining its initial orbit. Lastly, the fourth mission (0.D.5) is a multi-agent scenario in which four satellites learn to distribute themselves evenly along a GEO orbit.
These missions capture key aspects of both orbital dynamics and reinforcement learning (RL). Specifically, in terms of realism, we model various bodies, including drifters and satellites, with distinct characteristics such as radius, dry and fuel mass, and propulsion systems featuring different specific impulses and thrust magnitudes. From an orbital dynamics perspective, we train agents using simplified dynamics and evaluate their learned policies in more complex environments to assess generalization, with real perturbations present. Additionally, we leverage multiple coordinate systems and OrbitZoo’s built-in methods to facilitate efficient implementations across all missions.

### v2

@@ Appendix B Orbital Mechanics
This section provides an overview of key orbital mechanics concepts to support a clearer understanding of the concepts presented throughout the paper. We begin by explaining the most commonly used coordinate systems, then discuss our approach to thrust representation, and conclude with how body propagation is usually managed.
B.1 Coordinate Systems
@@
Figure 6: Representation of the Cartesian position ($r$) and velocity ($\dot{r}$) of a small body with mass $m$ orbiting a large central body with mass $M$.
Keplerian elements, as Figure 7 shows, describe the orbital motion using five parameters: the semi-major axis $(a)$, eccentricity $(e)$, inclination $(i)$, argument of perigee $(\omega)$ and longitude of the ascending node $(\Omega)$. Within this orbit, the anomaly (which can be represented in three ways, as seen in Figure 8) is an angle that indicates the current body position within that orbit, measuring from the periapsis. The mean anomaly is commonly used due to its linear evolution over time. When there are no perturbative forces (including maneuvers) that change the inclination, all points of the orbit are included in a two-dimensional plane, called the orbital plane. In these cases, Cartesian positions and velocities vectors are always within this plane.
Figure 7: Representation of Keplerian Orbital Elements. Here, it is assumed that the orbit is perfectly circular ($e=0$), and consequently, the semi-major axis ($a$) corresponds to the Euclidean distance to the primary focus ($M$).
Figure 8: Types of anomaly. A body $S$ orbits a central body $C$ in an elliptical orbit with semi-major axis $a$. The true anomaly ($\theta$) is the actual angular position of $S$ measured from the periapsis (closest point to $C$). The mean anomaly ($M$) is the angle assuming uniform motion over the same orbital period (represented by body $S^{\prime}$). Eccentric anomaly ($E$) is an intermediate angle used to relate $M$ and $\theta$, representing the position of $S$ on the circular orbit ($Q$). Despite evolving at different rates, all types of anomaly reach periapsis (0 rad) and apoapsis ($\pi$ rad) at exactly the same moment.
These elements directly relate to the orbital geometry and are particularly useful for characterizing two-body motion. However, $\omega$ and $\Omega$ can become undefined or ambiguous in special cases, such as circular or equatorial orbits. These scenarios, referred to as singularities, pose significant challenges for numerical integration and optimization problems. Equinoctial elements are therefore preferred for RL, as they avoid these singularities by using five parameters:
$a=a\quad e_{x}=e\cos(\omega+\Omega)\quad e_{y}=e\sin(\omega+\Omega)\quad h_{x}=\tan(i/2)\cos(\Omega)\quad h_{y}=\tan(i/2)\sin(\Omega),$
where $e_{x}$ and $e_{y}$ represent components of the eccentricity vector, and $h_{x}$ and $h_{y}$ describe the inclination vector.
B.2 Thrust Representation
@@
where $\times$ represents the cross product and $\|\cdot\|$ is the L2 norm. This coordinate system places the spacecraft at the origin, with $\hat{R}$ pointing towards the central body, $\hat{W}$ pointing perpendicular to the orbital plane, and $\hat{S}$ pointing perpendicular to $\hat{W}$ and $\hat{R}$ (approximately the same direction as $\dot{r}$).
@@
$).\mathbf{T}_{\text{RSW}}=T(\cos{\theta}\hat{S}+\sin{\theta}(\cos{\phi}\hat{R}+\sin{\phi}\hat{W}$)). (2)
@@ B.3 Propagation
$\mathbf{F}=-\frac{GMm}{\|r\|^{2}}\hat{r}$, (5)
where $G$ is the gravitational constant and $\hat{r}$ is the unit vector of position $r$. Using Newton’s second law, the resulting acceleration of the smaller body is expressed as:
@@
Although this model forms the foundation of orbital mechanics and is commonly used in RL, analytical models like SGP incorporate additional perturbative forces – such as atmospheric drag, SRP, and Earth’s oblateness – to provide a fast and relatively accurate method of propagating bodies.
Numerical propagators provide a powerful alternative to SGP by solving the equations of motion through numerical integration, enabling precise modeling of complex orbital dynamics. Unlike SGP, numerical propagators can incorporate a wide range of perturbative forces ($a_{\text{env}}$), including higher-order gravitational harmonics, third-body effects, atmospheric drag, and SRP. This flexibility allows them to handle scenarios requiring high precision, such as long-term orbit predictions and mission-critical maneuvers.
@@
$f(t,s)=\frac{d}{dt}s=\frac{d}{dt}\begin{bmatrix}r\\ \dot{r}\\ m\end{bmatrix}=\begin{bmatrix}\dot{r}\\ \ddot{r}\\ \dot{m}\end{bmatrix}=\begin{bmatrix}\dot{r}\\ -\frac{\mu_{E}}{\|r\|^{2}}\hat{r}+\frac{\mathbf{T}}{m}+a_{\text{env}}\\ -\frac{\|\mathbf{T}\|}{I_{\text{sp}}g}\end{bmatrix}$. (7)
@@
The time dependence of $f(s,t)$ in equation 7 enables the modeling of forces that are active only at specific times, such as environmental accelerations $a_{\text{env}}$ or thrust maneuvers.
To propagate the uncertainty, one can use the state transition matrix (STM) to analytically approximate the expected uncertainty found in Monte Carlo simulations.
Assuming we are only interested in the uncertainty of the position $r$ and velocity $\dot{r}$, the state can be simplified to $s=(r,\dot{r})$, therefore representing the dynamics as $f(t,s)=(\dot{r},\ddot{r})$. By calculating the Jacobian $A\in\mathbb{R}^{6\times 6}$, which linearizes the dynamics around $t_{0}$:
$A=\frac{\partial f}{\partial s}=\begin{bmatrix}\frac{\partial\dot{r}}{\partial r}&\frac{\partial\dot{r}}{\partial\dot{r}}\\ \frac{\partial\ddot{r}}{\partial r}&\frac{\partial\ddot{r}}{\partial\dot{r}}\end{bmatrix}=\begin{bmatrix}0&I\\ \frac{\partial\ddot{r}}{\partial r}&\frac{\partial\ddot{r}}{\partial\dot{r}}\end{bmatrix}$, (9)
it is possible to approximate the STM ($\Phi\in\mathbb{R}^{6\times 6}$) as:
$\Phi=I+A\Delta t.$ (10)
By knowing the position and velocity covariance matrix at $t_{0}$ ($\Sigma_{0}\in\mathbb{R}^{6\times 6}$), the propagated covariance ($\Sigma_{\Delta t}$) becomes:
$\Sigma_{\Delta t}=\Phi\Sigma_{0}\Phi^{T}.$ (11)
Appendix C Formal Model Definitions
@@
C.1 Double Deep Q-Network (DDQN)
Double Deep Q-Network (DDQN) [29] is an improvement over the standard Deep Q-Network (DQN) [51] algorithm used in RL for environments with discrete action spaces. Although powerful, DQN suffers from overestimation bias (the tendency to overestimate action values due to using the same network for both action selection and action evaluation). DDQN addresses this by decoupling these two roles using two networks: an online network with parameters $\theta$ used for selecting actions, $Q_{\theta}(s_{t},a_{t})$, and a target network with parameters $\theta^{-}$ used for evaluating the value of the selected actions without constant deviations, $Q_{\theta^{-}}(s_{t},a_{t})$.
After an agent interacts with the environment and stores several experience tuples $(s_{t},a_{t},r_{t},s_{t+1})$ in a buffer, DDQN samples a batch $B$ of experiences from the buffer and minimizes the mean squared error (MSE) between the current Q-value, $Q_{\theta}(s_{t},a_{t})$, and the estimated real Q-value, $Q^{*}_{\theta}(s_{t},a_{t})$:
$\min_{\theta}\frac{1}{|B|}\sum_{(s_{t},a_{t},r_{t},s_{t+1})\sim B}(\underbrace{r_{t}+\gamma Q_{\theta^{-}}(s_{t+1},\pi_{\theta}(s_{t+1}))}_{Q^{*}_{\theta}(s_{t},a_{t})}-Q_{\theta}(s_{t},a_{t}))^{2},$ (12)
where $|B|$ is the size of the batch, $0\leq\gamma<1$ is the discount factor and $\pi_{\theta}(s_{t+1})=\max_{a}Q_{\theta}(s_{t+1},a)$ corresponds to the best action according to the current parameters $\theta$. Periodically, the target parameters $\theta^{-}$ are updated with the parameters of the online network, $\theta$.
Exploration in DQN is guided by the epsilon-greedy strategy, where at each time step, the agent selects a random action with probability $\epsilon\in(0,1)$, and with probability $1-\epsilon$, it chooses the action with the highest estimated Q-value. This probability $\epsilon$ typically starts relatively high to encourage exploration early in training and gradually decreases over time, allowing the agent to increasingly exploit the strategies it has learned.
C.2 Deep Deterministic Policy Gradient (DDPG)
DDPG [43] is an off-policy RL algorithm that contains an actor and a critic network, as it stores experiences in a buffer gathered in interactions with old policies, similar to DQN. The actor network with parameters $\theta$ receives a state and produces an action (representing the policy, $\pi_{\theta}(a_{t}|s_{t})$), and the critic network with parameters $\phi$ receives a state and an action, and produces a single value (representing the Q-value function, $Q_{\phi}(s_{t},a_{t})$). Similarly to DQN, DDPG also contains a target actor with parameters $\theta^{-}$ and a target critic with parameters $\theta^{-}$.
For optimization, both networks are dependent from the value that the other outputs. While the actor network tries to maximize the expected value produced by the critic (equation 13), the latter tries to approximate the Q-values to the true estimated ones by a MSE loss function (equation 14):
$\max_{\theta}E\left[Q_{\phi}(s_{t},\pi_{\theta}(s_{t})\right]$ (13)
$\min_{\phi}\frac{1}{|B|}\sum_{(s_{t},a_{t},r_{t},s_{t+1})\sim B}(\underbrace{r_{t}+\gamma Q_{\phi^{-}}(s_{t+1},\pi_{\theta^{-}}(s_{t+1}))}_{Q^{*}_{\phi}(s_{t},a_{t})}-Q_{\phi}(s_{t},a_{t}))^{2}.$ (14)
The target network can be updated using the soft update method, rather than directly copying the online network. This method uses a parameter $\tau\in(0,1)$ to keep a percentage of the old targets, preventing large, abrupt changes in target values:
$\theta^{-}\leftarrow\tau\cdot\theta+(1-\tau)\cdot\theta^{-}.$ (15)
Exploration in DDPG can be achieved using the Ornstein-Uhlenbeck process [76], which introduces temporally correlated noise $x_{t}$ to the actions produced by the actor network. This noise evolves over time and is updated at each time step according to:
$x_{t+1}=x_{t}+\theta(\mu-x_{t})\Delta t+\sigma\sqrt{\Delta t}\cdot\mathcal{N}(0,1),$ (16)
where $\theta$ determines how strongly the noise is pulled back toward the mean $\mu$, $\sigma$ controls the magnitude of the noise, and $\Delta t$ is a small time increment used to discretize the continuous process.
C.3 Proximal Policy Optimization (PPO)
The actor loss in PPO [66] is defined based on the ratio $r_{t}(\theta)$, which measures the probability of taking an action $a_{t}$ under the current policy compared to the old policy:
@@ C.3 Proximal Policy Optimization (PPO)
Since the actor network in PPO outputs the expected action the agent should take given a state, exploration is naturally handled by sampling from a probability distribution. Specifically, an action is sampled from a normal distribution centered at the policy output with a given standard deviation $\sigma$:
$a_{t}\sim\mathcal{N}(\pi_{\theta}(a_{t}\mid s_{t}),\sigma).$ (21)
The standard deviation $\sigma$ typically starts high to encourage exploration in the early stages of training and is gradually decreased over time to allow the agent to exploit the strategies it has learned.
C.4 Generalized Advantage Estimation (GAE)
Generalized Advantage Estimation (GAE) [65] provides a flexible way to compute the advantage by balancing bias and variance through the $\lambda$ parameter:
@@ C.4 Generalized Advantage Estimation (GAE)
C.5 Federated Learning (FL)
In MARL environments, there are several approaches to collaboratively train agents. A common method is independent learning, where each agent is trained separately and treats other agents as part of the environment’s dynamics. In this setting, collaboration arises indirectly through a shared reward function, which must be identical for all agents to ensure coordinated behavior.
However, Federated Learning (FL) offers a more stable and scalable alternative by enabling agents to share knowledge during training while keeping their local data private. FL is a machine learning paradigm in which multiple entities (e.g., agents or devices) collaboratively train a global model without sharing raw data. Instead, they exchange model parameters or gradients, which are aggregated to improve a shared model.
Applied to RL, FL can be used to exchange learned parameters between agents rather than raw experience tuples. This is particularly beneficial in POMDP environments, where each agent has only a limited view of the full state. By federating value functions or policy networks, agents can better estimate the value of a given state $s_{t}$, leading to more stable and data-efficient learning while maintaining a level of privacy and decentralization.
A specific case of FL is horizontal federated learning (HFL) [56], which occurs when all agents have the same observation and action spaces. In such scenarios, multiple identical agents operating in parallel can be treated as instances of a single agent in multiple environments. This setting allows for straightforward parameter sharing and can be seen as data augmentation across parallel environments, enhancing generalization and learning speed.
One common HFL algorithm that can be used for recent RL algorithms (e.g. DDPG and PPO) is Federated Averaging (FedAvg) [48], by training a shared critic through a central server (e.g., ground station). In an environment with $N$ agents, the server starts by sending the current global parameters ($w_{g}$) to every agent: $w_{i}\leftarrow w_{g},\forall i\in\{1,...,N\}$ (where $w_{i}$ are the local parameters of agent $i$). Then, agents interact with the environment for an arbitrary number of steps, gathering a buffer of experiences $B_{i}$, which are then used to locally optimize the weights $w_{i}$ by an RL algorithm. After training, these weights are gathered by the server and averaged:
$w_{g}=\sum_{i=1}^{N}\frac{|B_{i}|}{\sum_{k=1}^{N}|B_{k}|}w_{i},$ (23)
where $|B|$ is the number of experiences in buffer $B$. Finally, the process is repeated by sending these new global parameters to each agent. When we assume that the experiences gathered by each agent are independent identically distributed (IID), FedAvg converges.
Appendix D Computational Performance
@@
OrbitZoo imposes no strict limit on the number of bodies in a system, with scalability constrained only by the available hardware (D.1). However, it is essential to assess how increasing the number of bodies and introducing complex dynamics impact simulation speed and parallelization efficiency. These aspects are examined in the following subsections (D.2 and D.3).
Although not yet implemented, OrbitZoo experiments could benefit from libraries mentioned in 3.3 to accelerate training.
D.1 Hardware Specifications
@@ D.2 Scalability
To address the first question, we evaluate system performance when propagating 1, 5, 10, 100, 1000, and 10000 bodies. For the second question, we analyze simulation speed starting with a simple Newtonian gravity model assuming a perfectly spherical central body. We then introduce a more complex gravity model using spherical harmonics (HolmesFeatherstone) and further incorporate perturbative forces, including third-body effects (Sun and Moon), SRP, and drag.
@@
Figure 10: Time per step (in seconds) by number of bodies and active forces. Each point represents an average of 100 propagation steps of 10 seconds. Shaded regions indicate $\pm 1$ standard deviation.
D.3 Parallelization
@@
As shown in Figure 4, OrbitZoo integrates with two key external components: PettingZoo and Orekit. PettingZoo provides a framework for sequential or parallelized step computations, while Orekit handles propagation of orbital dynamics within each step. Since orbital MARL missions assume all agents act simultaneously rather than in turns, OrbitZoo propagates systems in a parallel manner, so it should naturally take advantage of available tools for faster computations.
@@
OrbitZoo leverages PettingZoo’s parallelization by implementing the ParallelEnv class, and Orekit’s parallelization by implementing the PropagatorsParallelizer class (a way to simultaneously propagate several bodies). To assess the impact of this choice, we compared simulation times between parallel and sequential step computations (as it was presented in Figure 10).
Figure 11 shows that when using few active forces, the sequential propagation is faster than the parallel, even for a large number of bodies. However, when systems become more computational demanding – with many perturbative forces – the parallel mode becomes the fastest. Given that a single RL episode can consist of hundreds or thousands of steps, this improvement significantly enhances simulation efficiency, where using the parallel mode can decrease a single episode by several minutes.
Because of these results, OrbitZoo allows the user to change the computation mode upon environment creation (boolean value) according to its needs (through the JSON input file).
@@
Figure 11: Time per step (in seconds) when using sequential and parallel computations, by the number of bodies and active forces. Each point represents an average 100 propagation steps of 10 seconds. Shaded regions indicate $\pm 1$ standard deviation.
D.4 Challenges and Future Improvements
As mentioned earlier, each propagator is currently treated independently, with duplicated instances for similar bodies. Future work regarding scalability may be related to automatically eliminating this redundancy for body propagators, as forces may be similar for different bodies, therefore benifitting from using the same force instances. This effectively reduces memory usage without the need to manually change these settings.
As real-world satellite constellations continue to grow in size and complexity, further research into scalability becomes increasingly important. Ensuring that RL-based control systems can scale effectively with the number of satellites, while maintaining performance and reliability, is a critical step toward practical deployment in real space missions.
Appendix E Experiments
To evaluate OrbitZoo’s capabilities, we designed a series of missions utilizing algorithms ranging from foundational to state-of-the-art (as discussed in Section E.1). The first mission (Section E.2) compares the OrbitZoo environment to D. Kolosa’s implementation for a general orbit change maneuver using DDPG, highlighting how the tools and flexibility provided by OrbitZoo can support the development of both existing and novel, more complex missions – without the technicalities typically present in orbital dynamics.
Next, two missions (Sections E.3 and E.4) focus on orbital transfers (OTs) using PPO, where agents learn strategies to reach static and dynamic targets, respectively, while considering real-world factors such as perturbative forces and mass loss due to fuel consumption.
We then present a collision avoidance maneuver (CAM) mission (Section E.5), in which a satellite learns to minimize the probability of collision with debris in the days leading up to a predicted close approach, while maintaining its initial orbit. This showcases how OrbitZoo can also be employed in missions where uncertainty plays a central role. In this mission, we directly compare the performance of strategies learned using DDQN and PPO, highlighting how discrete RL algorithms perform relative to continuous ones in CAM missions.
Finally, we introduce a multi-agent reinforcement learning (MARL) mission (Section E.6) set in Geostationary Earth Orbit (GEO), where four satellites learn to distribute themselves evenly along the orbit without losing altitude. For this scenario, we compare independent learning and federated learning approaches using the PPO algorithm, demonstrating how knowledge sharing among agents in cooperative settings can be beneficial.